### v1

# Do Androids Know They’re Only Dreaming of Electric Sheep?

Abstract
We design probes trained on the internal representations of a transformer language model that are predictive of its hallucinatory behavior on in-context generation tasks. To facilitate this detection, we create a span-annotated dataset of organic and synthetic hallucinations over several tasks. We find that probes trained on the force-decoded states of synthetic hallucinations are generally ecologically invalid in organic hallucination detection. Furthermore, hidden state information about hallucination appears to be task and distribution-dependent. Intrinsic and extrinsic hallucination saliency varies across layers, hidden state types, and tasks; notably, extrinsic hallucinations tend to be more salient in a transformer’s internal representations. Outperforming multiple contemporary baselines, we show that probing is a feasible and efficient alternative to language model hallucination evaluation when model states are available.
1 Introduction
Do language models know when they’re hallucinating? [^1] That is, do they know that their electric sheep are ungrounded? Detecting hallucinations in grounded generation is commonly framed as a textual entailment problem. Prior work has largely focused on creating secondary detection models trained on and applied to surface text Ji et al. (2023). However, these secondary models are often of comparable complexity to the underlying language model that generated the error and ignore the information that was already computed during generation. In this work, we explore the degree to which probes on a decoder-only transformer language model’s feed-forward and self-attention states can detect hallucinations in a variety of grounded generation tasks.
Figure 1: Prediction of hallucination in a generated response by probing the hidden states of a transformer during decoding. The true annotated span of hallucination within the response is boxed in white; rows represent probes trained to detect hallucination from layers within a transformer’s hidden states, with columns representing their prediction for each token in a generated response.
In particular, as the language model generates a response, we probe its encoded tokens to determine whether the model has just begun to hallucinate or will eventually be found to have hallucinated. It is plausible that these encodings are “aware” of whether generation is retrieving information that is grounded in the prompt versus filling in plausible information that is stored in the model’s weights. Related work has studied how transformers retrieve relevant memories as they encode text Meng et al. (2022, 2023) and suggests that feed-forward layers act as associative memory stores Geva et al. (2021). A probe for hallucination detection must consider whether subsequent generated text is entailed by information in the prompt, given the retrieved memories, or is merely one plausible extrapolation from the prompt (see also Monea et al., 2023).
We develop three probes of increasing complexity across three grounded generation tasks: abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. For each task, we collect hallucinations in two ways: (1) from sampled responses, where we generate outputs from a large language model (LLM) conditioned on the inputs, or (2) by editing reference inputs or outputs to create discrepancies. We refer to these as organic and synthetic data, respectively. In both cases, we have human annotators mark hallucinations in the outputs. Method (2) produces hallucination annotations at a higher rate, though we find that the utility of these synthetic examples is lower as they do not come from the test distribution.
We summarize our contributions as follows:
  • We produce a high-quality dataset of more than 15k utterances with hallucination annotations for organic and synthetic output texts across three grounded generation tasks.
  • We propose three probe architectures for detecting hallucinations and demonstrate improvements over multiple contemporary baselines in hallucination detection.
  • We analyze how probe accuracy is affected by annotation type (synthetic/organic), hallucination type (extrinsic/intrinsic), model size, and which part of the encoding is probed.
2 Related Work
Definitions.
We center our study of hallucinations in the setting of in-context generation Lewis et al. (2020) where grounding knowledge sources are provided within the prompt itself. This is in contrast to settings where a language model is expected to generate a response entirely from its world knowledge (i.e., grounded in the model’s training data) Azaria and Mitchell (2023).
Though the term has often appeared in reference to factuality Gabriel et al. (2021), hallucinations are not necessarily factually incorrect. In their survey, Ji et al. (2023) classify hallucinations as intrinsic, where generated responses directly contradict the knowledge sources, or extrinsic, where generated responses are neither entailed nor contradicted by the sources. We follow this taxonomy. For example, given the following task prompt for abstractive summarization,
…Deckard is now able to buy his wife Iran an authentic Nubian goat—from Animal Row, in Los Angeles—with his commission …TL;DR:
A generated summary sentence of “Deckard can now purchase an animal for his wife” is faithful, “Deckard can purchase a goat in Iran” is an intrinsic hallucination, and “Sheep are usually cuter and fluffier than goats” is an extrinsic hallucination. While models may leverage latent knowledge to inform their responses, such as recognizing that a goat is indeed an animal, directly stating such knowledge in a response—i.e. only generating “Goats are animals” in the example above—is often considered an extrinsic hallucination as it is neither entailed nor contradicted by a given knowledge source. Thus, non-hallucinatory generation in this setting involves balancing the language model’s retrieval and manipulation of knowledge, both from the given context and from its own stored weights.
Task History Knowledge Instruction Response
Abstractive …dialed down that initial TL;DR: The National Weather
Summarization forecast to predict 13 to 19 Service forecast called
named storms, including six for 13 to 19 named
to nine hurricanes… storms including six
to 11 hurricanes.
KGDG …B: ok how can I get a Three consecutive strikes Response that uses A: Indeed, you have to
turkey I think that may are known as a "turkey." the information in get four consecutive
be hard to get though knowledge: strikes to get one.
Data-to-Text name[The Wrestlers], Restaurant Located by the river-
priceRange[high] Description: side, The Wrestlers is
a restaurant with high
prices, and near Raja
Indian Cuisine.
Table 1: From top to bottom, examples of organic hallucinations in model responses for abstractive summarization, knowledge-grounded dialogue generation (KGDG), and data-to-text tasks. Minimum annotated hallucination spans are highlighted in cyan for intrinsic hallucinations and green for extrinsic hallucinations.
Automated Diagnosis.
Lexical metrics commonly used in the evaluation of NLG fail to correlate with human judgments of faithfulness Kryscinski et al. (2019). More useful are NLI approaches that either directly measure the degree of entailment between a generated response and a given source text Hanselowski et al. (2018); Kryscinski et al. (2020); Goyal and Durrett (2020), or else break down texts into single-sentence claims that are easily verifiable Min et al. (2023); Semnani et al. (2023); these have been successfully used for hallucination detection Laban et al. (2022); Bishop et al. (2023). Similarly, approaches that leverage question-answer models to test if claims made by a response can be supported by the source text, and vice versa Wang et al. (2020); Durmus et al. (2020); Scialom et al. (2021), try to more explicitly identify the grounding sources of given generations in this task. Under the intuition that a model is usually not confident when it is hallucinating, hallucinatory behavior can also be reflected in metrics that measure generation uncertainty or importance Malinin and Gales (2021); Guerreiro et al. (2023); Ren et al. (2023); Ngu et al. (2023); Madsen et al. (2023), relative token contributions Xu et al. (2023), and differences between the attention matrices and decoder stability Lee et al. (2018) during generation.
Predicting Transformer Behavior.
In small transformer language models, it is possible to ascribe behaviors like in-context learning to specific layers and modules Elhage et al. (2021). This becomes harder in larger language models, but Meng et al. (2022) provide evidence that the feed-forward modules function as key-value stores of factual associations and that it is possible to modify weights in these modules to edit facts associated with entities. Our probes, to be successful, must determine whether the recalled facts were deployed in accordance with the rules of grounding (i.e., to paraphrase facts expressed in the prompt, rather than to replace them or to add details that are merely probable). Transformer internal states have also been shown to be predictive of the likelihood that a model will answer a question correctly in the setting of closed-book QA Mielke et al. (2022) and in the prediction of the truthfulness of a statement Azaria and Mitchell (2023), though these internal representations of truthfulness can disagree systematically with measures of model confidence Liu et al. (2023). These works measure hidden states in the absence of supporting evidence (i.e. no passages are retrieved for either task), making them effectively probes of knowledge stored in model weights and not the manipulation of novel or contingent information provided in a prompt. Monea et al. (2023) find that there are separate computational processes in transformer LLMs devoted to recall of associations from memory vs grounding data, and we show that we can detect whether associations from grounding data are used correctly across several distinct generation tasks. Slobodkin et al. (2023) train a probe to detect the (un)answerability of a question given a passage, which is similar in some respects to our task in Section 4; however, we detect token-level hallucination during answer generation, which could happen even in theoretically answerable questions.
Response Original Knowledge Changed Knowledge
Yes, Elvis was also one of Elvis Presley is regarded as one of Elvis Presley is regarded as one
the most famous musicians the most significant cultural icons of the least well-known music-
of the 20th century. of the 20th century. -ians of the 20th century.
Close to The Sorrento you name[Clowns], eatType[pub] name[Clowns],
can find Clowns pub. near[The Sorrento] near[The Sorrento]
Table 2: Examples of changed knowledge statements for our datasets of synthetic hallucinations. Changes made to Conv-FEVER (top) and E2E (bottom) are shown in bold. The changed statements make the top response an intrinsic predicate error and the bottom response an extrinsic entity error (highlighted as in Table˜1).
3 Grounded Generation Tasks
We test hallucination probes for autoregressive grounded generation in three distinct tasks:  abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. We use the same high-level prompt structure across all tasks to organically generate model responses to tasks and to probe for hallucination. Each prompt contains an optional interaction history, a knowledge source, and an instruction resulting in the conditionally generated response. Task prompts and examples of organic hallucinations are shown in Table˜1.
3.1 Organic Hallucinations
We use llama-2-13b Touvron et al. (2023) as the response model to generate outputs for each task. To ensure that a non-negligible number of both hallucinations and grounded responses were present in the generations for each task, each model response is generated using top-k random sampling Fan et al. (2018) with $k=2$. The number of few-shot in-context learning examples for each task was chosen through a pilot exploration of several generation strategies, optimizing for the presence of a decent balance of hallucinations and grounded responses in all tasks, as determined via manual inspection by the authors. Full prompt templates are shown in Appendix˜C.
Abstractive Summarization.
The CNN / DailyMail (CDM) dataset Hermann et al. (2015) is a standard benchmark in the setting of abstractive summarization Lin and Ng (2019). It consists of a set of news articles paired with reference summaries in the form of human-written lists of bullet points that were originally published alongside the articles. For automatic summarization, we follow the GPT-2 paper Radford et al. (2019) and prompt our model to generate abstractive summaries of news articles by appending TL;DR: after the article content (a 0-shot prompt), [^2] In Internet posts that appear in LLM pretraining data, this string often marks the start of a summary immediately following the main content of the post. Prior work Radford et al. (2019) has found that appending this string after texts in prompts to autoregressive LLMs is capable of producing abstractive summaries of the former. taking the first three generated sentences as the summary.
Knowledge-Grounded Dialogue Generation.
For a given conversation history and a provided knowledge sentence obtained via Wizard-Of-Wikipedia, the Conv-FEVER Santhanam et al. (2021) task (CF) asks a respondent to create a next-turn response to the conversation that is grounded in the provided knowledge sentence. Here, we prompt our response model to generate grounded responses under 2-shot in-context learning with example grounded responses.
Data-to-Text Generation.
The E2E task Novikova et al. (2017) asks respondents to write restaurant descriptions that are congruent with a provided set of attributes. Here, we prompt our response model to perform this task with 10-shot in-context learning. [^3] As the original E2E task had omitted or missing information in up to 40% of its data, we use the cleaned version by Dušek et al. (2019). As even the cleaned version had instances of omitted or missing information, 10 in-context examples were manually selected and manually verified for groundedness.
3.2 Synthetic Hallucinations
Recent work Ma et al. (2023); Monea et al. (2023) has carefully created synthetic ungrounded datasets by hand to evaluate metrics used to measure unfaithfulness and models used to detect it. The BUMP dataset Ma et al. (2023) is a version of CDM in which reference summaries have been manually edited to be ungrounded. It includes both freestyle edits and systematic coverage of 4 specific types of faithfulness errors: (1) predicate errors, where the predicate in a generation is inconsistent with the source knowledge; (2) entity errors, where the subject or object of a predicate is inconsistent; (3) circumstance errors, or errors in the time, duration, or location of an event of the predicate; and (4) coreference errors. Types (1)–(3) may be intrinsic or extrinsic.
We use BUMP in our CDM experiments. For our Conv-FEVER and E2E experiments, we similarly created faithfulness errors in a uniformly sampled subset of these datasets. However, rather than changing an example’s reference response, we changed its provided knowledge sentence so that the existing reference response would be unfaithful to it. This simulates the common case of hallucination where a large language model generates statements that may be true, but which are not grounded in the source documents as required. For Conv-FEVER, we prompted ChatGPT to modify a given example’s provided knowledge sentence to introduce intrinsic and extrinsic errors of types (1)–(4) above; we then manually judged these attempts and edited the unsuccessful ones. [^4] In hindsight, this workflow proved inefficient; significant human editing and validation were needed. In this way, we created equal numbers of examples with intrinsic and extrinsic predicate, entity, circumstance, and coreference errors, as well as freestyle errors.
Similarly, for E2E, we modified a random subset of the provided grounding attributes. Specifically, for an example with $n$ attributes, we drew $k$ uniformly from [1,n1][1,n-1], uniformly selected one of the subsets of size $k$, and then flipped a fair coin to decide whether to remove or perturb—by randomly sampling an alternative attribute from the entire dataset—these $k$ attributes. Finally, we manually verified that the reference response was unfaithful with respect to the modified attributes; if not, we further edited the attributes. Examples of changed knowledge statements are shown in Table˜2; statistics of our synthetic datasets are shown in Table˜3.
Task N-o H-o N-s H-s
CDM 3399 23% 1778555Obtained via BUMP Ma et al. (2023). 50%
CF 924 64% 2846 50%
E2E 4975 57% 9950 50%
Table 3: Dataset statistics. N-o is the number of annotated organic generations, and H-o is the percentage of organic generations with at least one hallucination. N-s and H-s refer to the same for synthetic responses. Note that grounded synthetic responses were not generated by a model but come from gold reference data.
3.3 Annotation
We ask annotators to label the minimal hallucinatory spans in each response, if any. These are the smallest text spans that would need to be modified to transform an ungrounded response into a grounded one.
A team of 17 annotators completed a pilot annotation exercise of 50 annotations per task. These pilot annotations were checked and verified by the authors to pinpoint common sources of error and disagreement. This feedback was integrated into a revised set of annotation guidelines, shown in Appendix˜A, and the full dataset was then reannotated by the same team. Each example was annotated once; example annotations are shown in Table˜1. The number of annotated examples is shown in Table˜3. Note that for each dataset, over one-fifth of all annotated responses were annotated as containing at least one hallucinated span (columns H-o and H-s).
4 Probing
Figure 2: General probe structure. Response tokens that were annotated as hallucinatory are shown in red. For each internal representation layer, and each type of representation (attention output and feed-forward output), a classifier is separately trained on the attention head out-projection and feed-forward hidden states for all representation layers (or just the final token) of a response from a prompted language model. Depending on the probe, these classifiers are then ensembled to make a final prediction of whether each individual token was, or otherwise if the response contained, a hallucination.
Probes Alain and Bengio (2017) are tools used to analyze a neural network’s internal representations. Often taking the form of linear classifiers, they are applied to the internal representations of a network and are trained to discriminate between different types of inputs or outputs. Here, we are interested in detecting hallucinatory outputs during an LLM’s generation steps when conditioned on a prompt that requires a response grounded n the prompt.
4.1 Design
Let $u$ be a prompt as described at the start of Section 3. To determine whether a response $x$ to a prompt $u$ contains hallucinations, we will probe the hidden states of the transformer language model (i.e., decoder) as it generates $x$. Note that these hidden states depend on $u$ and on the prefix of $x$ that has been generated so far and that they may be reconstructed from the string pair $(u,x)$ via forced decoding. A single transformer layer actually consists of two sublayers: the first transforms a token’s embedding by adding a vector predicted by attention, and the second transforms it again by adding a vector predicted by a feed-forward network. We train a separate probe for each of the $2L$ sublayers, where $L$ is the number of layers.
Linear Probe.
This probe is a linear classifier using a single hidden state vector (output by some sublayer) as input. Let $x_{i}$ be the $i^{\textrm{th}}$ response token and let $\mathbf{h}_{i}\in\mathbb{R}^{n}$ be the corresponding LLM hidden state at a fixed layer and component. We write $y_{i}\in\{0,1\}$ to indicate whether $x_{i}$ falls within a minimum hallucination span. The linear probe predicts $y_{i}$ by $p(y_{i}=1\mid u,x_{\leq i})=\sigma(\mathbf{w}^{\intercal}\mathbf{h}_{i}+b)$, where $\mathbf{w}\in\mathbb{R}^{n}$ are learned weights, $b\in\mathbb{R}$ is a learned bias, and $\sigma$ is the sigmoid function.
Attention-Pooling Probe.
In order to directly consider the previous hidden states of the response as well, we also experiment with an attention-pooling probe over all hidden states so far,
$p(y_{i}=1\mid u,x_{\leq i})=\sigma(\mathbf{w}^{\intercal}\mathbf{\bar{h}}_{i})$
$\mathbf{\bar{h}}_{i}=\sum_{j=1}^{i}\alpha_{i,j}\mathbf{h}_{j},\;\;\alpha_{i,j}=\frac{\exp\left(\mathbf{q}^{\intercal}\mathbf{h}_{j}\right)}{\sum_{k=1}^{i}\exp\left(\mathbf{q}^{\intercal}\mathbf{h}_{k}\right)}$
where $\mathbf{q}\in\mathbb{R}^{n}$ is a learned query vector.
Ensemble Probe.
Our final approach is a single feed-forward layer that combines the predictions from our $2L$ separately trained probes (either the linear probes or the attention-pooling probes):
$p(y_{i}\mid u,x_{\leq i})$ $=\sigma\left(\sum_{\mathclap{\begin{subarray}{c}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;l\in\{1,\ldots,L\}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;m\in\{\textrm{attention},\textrm{feed-forward}\}\end{subarray}}}\beta_{l,m}\cdot p_{l,m}(y_{i}\mid u,x_{\leq i})\right)$
where $\beta_{l,m}$ are learned weights. The individual probes are frozen before $\beta_{l,m}$ parameters are fit.
4.2 Training Objectives
Token-level.
To train a probe to predict token-level hallucination indicators $y_{i}$, we minimize negative log-likelihood on the annotations in our dataset $\mathcal{D}$:
$\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}\sum_{i=1}^{|x|}-\log p(y_{i}\mid u,x_{\leq i};\theta)$
Response-level.
We also try training a classifier to predict whether $x$ contains any hallucinations at all; denote this indicator as $y=\bigvee_{i=1}^{|x|}y_{i}\in\{0,1\}$. For this, we simply use an attention-pooling probe over the entire response $x$. To fit the probe parameters, $\theta$, we again minimize negative log-likelihood:
$\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}-\log p(y\mid u,x_{\leq|x|};\theta)$
4.3 Metrics
Response-level Classification.
F1 is the harmonic mean of recall and precision. We denote the F1 score of a response-level classifier by F1-R.
Token-level Classification.
For token-level classifiers, it is possible to evaluate the F1 score of their per-token predictions; however, this would give more weight to annotated or predicted spans that are longer in length. Here, we are more interested in span-level F1: specifically, whether we predicted all of the annotated hallucinations (recall), and whether all of the predicted spans were annotated as actual hallucinations (precision). As requiring an exact match of the spans is too severe—Appendix˜B shows how our human annotators often disagreed on the exact boundaries of a hallucinatory span—we modify span-level precision and recall metrics to give partial credit, as follows.
Define a span to be a set of tokens, where tokens from different responses are considered distinct. Let $A$ be the set of all annotated spans in the reference corpus, and let $P$ be the set of all predicted spans on the same set of examples. We compute the recall $r$ as the average coverage of each reference span, i.e., the fraction of its tokens that are in some predicted span. Conversely, we compute the precision $p$ as the average coverage of each predicted span:
$r$ $=\frac{1}{|A|}\sum_{s\in A}\frac{\left|s\cap\left(\bigcup_{\hat{s}\in P}\hat{s}\right)\right|}{|s|}$
$p$ $=\frac{1}{|P|}\sum_{\hat{s}\in P}\frac{\left|\hat{s}\cap\left(\bigcup_{s\in A}s\right)\right|}{|\hat{s}|}$
where $\cup$, $\cap$, and $|\cdot|$ denote set union, intersection, and cardinality respectively. We define F1-Sp (standing for "span, partial credit") as the harmonic mean of $p$ and $r$. Note that F1-Sp micro-averages over examples, ignoring the boundaries between them. Thus, in contrast to F1-R, examples with more spans have more influence on the final F1-Sp metric achieved.
5 Experiments
Unless otherwise indicated, all evaluations are performed on organic data only, even if their probes were trained on synthetic data.
5.1 Probe Hyperparameters and Training
Single-layer hidden state probes, described in Section 4, are trained under float32 precision with unregularized log loss. A batch size of 100 and Adam optimization Kingma and Ba (2014) with a learning rate of 0.001 and $\beta_{1},\beta_{2}=(0.9,0.999)$ were chosen after a small grid search on hyperparameters. With 70/10/20% train, validation, and test splits, each classifier was trained using early stopping with 10-epoch patience. With the predictions of these single-layer probes as input, the ensemble classifier was trained with the same hyperparameters.
5.2 Response-Level Baselines
We compare the performance of our probes against a set of contemporary methods used in faithfulness evaluation, measuring test set hallucination classification F1. Specifically: (1) the model uncertainty-based metric of length-normalized sequence log-probability Seq-Logprob Guerreiro et al. (2023); (2) off-the-shelf synthetically-trained sentence-level claim classification FactCC Kryscinski et al. (2020); (3) linguistic feature-based dependency arc entailment DAE Goyal and Durrett (2020); (4) QA-based metrics FEQA Durmus et al. (2020); (5) QuestEval Scialom et al. (2021); (6) the NLI-based metric SummaC Laban et al. (2022); and ChatGPT classifiers Luo et al. (2023) (7) without and (8) with chain-of-thought prompting, with prompts shown in §˜C.1.
Response-Level Classification, F1-R
Method CDM CF E2E
OC 0.38 0.78 0.72
Seq-Logprob 0.44 0.81 0.75
FactCC 0.43 - -
DAE 0.37 - -
FEQA 0.35 - -
QuestEval 0.41 - -
SummaCZS 0.46 - -
SummaCConv 0.42 - -
ChatGPT 0.34 0.62 0.37
ChatGPT CoT 0.36 0.68 0.45
PoolingSL 0.43 0.86 0.86
PoolingE 0.45 0.88 0.87
PoolingSL-SYNTH 0.46 0.64 0.66
PoolingE-SYNTH 0.48 0.65 0.69
Human 0.71 0.87 0.87
Table 4: Response-level probe and baseline organic hallucination detection performance across tasks as measured via F1 scores (F1-R) achieved on the test set. SL denotes the best single-layer classifier (chosen on the validation set), E denotes the ensembled classifier, and SYNTH indicates that the classifier was trained on synthetic data and tested on organic hallucinations. The Human row measures the performance of a human annotator who is different from the reference annotator; this was measured on a subset of 50 examples that were triply annotated, averaging over all 6 (annotator, reference annotator) pairs.
Metrics (2)–(6) were not originally intended to detect hallucinations in individual responses, but rather to evaluate a collection of responses for the sake of system comparisons. To adapt to our setting, however, we threshold each of these continuous metrics to classify the individual responses, setting the threshold to maximize validation set F1, following Laban et al. (2022). For sentence-level optimized baselines such as FactCC, response-level evaluation is performed by logical OR-ing sentence-level baseline predictions (obtained after thresholding the raw scores). For CDM, the synthetic training dataset (from BUMP) is smaller than the organic training dataset, so the synth results are at a disadvantage. For CF and E2E, however, we generated a synthetic training dataset that was matched in size.
As a further reference baseline, we additionally report the performance achieved by randomly predicting a hallucination label with some probability p tuned to maximize F1 on the validation set, denoted here as Optimized Coin (OC).
Refer to captionRefer to captionRefer to caption
Figure 3: Grounding behavior saliency—as measured by probe response-level F1 on test set—of single layer probe classifiers trained on the attention-head out projections and feed-forward states of llama-2-13b, stratified across tasks and layers.
Span-level Classification, F1-Sp
Method CDM CF E2E
OC 0.09 0.58 0.15
LinearSL 0.24 0.65 0.46
LinearE 0.27 0.66 0.47
PoolingSL 0.32 0.68 0.46
PoolingE 0.34 0.68 0.47
Human 0.61 0.69 0.58
Table 5: Span-level probe hallucination detection performance, as measured via span-level partial match scores (F1-Sp). The Human row is computed as in Table˜4 and is relatively stable across tasks, unlike the systems.
5.3 Results
Response-level Results.
On two datasets, we were actually able to match the interannotator agreement rate (the “Human” row). On these datasets, the probes trained on organic hallucinations worked best (Table˜4). The exception was the CDM dataset, where the organic probes lost slightly to the synthetic probes, which in turn fell well short of human performance. On all datasets, we were able to outperform all baseline systems, showing that probing is a feasible (and efficient) alternative to evaluating LLM outputs when model states are available.
Relative to other baselines that also have some access to model internals—such as Seq-Logprob—we hypothesize that probing-based approaches perform better here due to being able to capture information about model uncertainty that is lost when reduced to the output token probability.
Span-Level Results.
These results appear in Table˜5. Again, the “Human” row serves as a ceiling on how well we can expect our automated system to perform on these datasets, given that even humans do not fully agree. We observe that our ensemble attention-pooling probe nearly meets this ceiling for Conv-FEVER but lags behind for CDM and E2E. Still, even this level of performance is potentially useful for avoiding hallucinations, as we discuss in §˜6.
Layers.
Evaluating the performance of trained single-layer probes (Figure˜3), we see that signals of hallucination are strongest in model middle layers during decoding. Probe performance shows a consistent and gradual increase from the zeroth layer up until layer 10 to 20, after which there is a consistent decline in hallucination classification performance for the subsequent layers.
Response-level Classification, F1-R
Base Model Size CDM CF E2E
7B 0.45 0.88 0.87
13B 0.45 0.88 0.87
70B 0.45 0.89 0.87
Span-level Classification, F1-Sp
Base Model Size CDM CF E2E
7B 0.65 0.77 0.46
13B 0.64 0.77 0.46
70B 0.64 0.76 0.46
Table 6: Response and span-level F1 hallucination detection performance, as measured by F1-R and F1-Sp, respectively, of the PoolingE ensemble probe across llama-2 response model sizes and test sets across tasks. For each task and model size, a probe is trained on labeled responses of the same task that were generated organically from the 13B model.
Hidden State Type.
Slight variations in grounding saliency—the strength of the signal of hallucination as able to be detected by a probe, measured by probe hallucination classification performance—are also present among hidden states of different types during decoding. As shown in Figure˜3, feed-forward probes reach peak performance at slightly deeper layers than their attention head out-projection probe counterparts and exceed them in single-layer hallucination detection. Training ensemble classifiers on only the hidden states of the same type at every layer, we see how feed-forward hidden states possess a slight but statistically significant greater saliency of model grounding behavior.
Model Size.
Training probes in the same manner on the force-decoded hidden states of responses on our dataset using llama-2-7b and llama-2-70b, we find no statistically significant difference (via Kolmogorov–Smirnov) in hallucination detection across the hidden states from base models of varying sizes (Table˜6). These results show that we can still detect a hallucination even when it was originally generated from a different model, and that even the lower-dimensional hidden states of the 7B model still contain enough information to train a successful probe.
Figure 4: Intrinsic and extrinsic hallucination saliency across layers for abstractive summarization as measured by response-level probe recall on hallucination types.
Hallucination Type.
Next, we test for the saliency of different forms of hallucination behavior in model internal states. Here, with annotator-labeled test set responses according to whether they contained intrinsic or extrinsic hallucinations, we evaluate trained probe performance in correctly detecting the presence of hallucinations for instances of each hallucination type. We observe in Figure˜4 that extrinsic hallucination behavior is more salient in the hidden states of shallower layers for both organic and synthetic hallucinations, with saliency noticeably peaking around layers 15 and 16, decreasing in deeper layers. Overall, extrinsic hallucination remained more salient in model hidden states than intrinsic hallucination.
Synthetic Hallucinations.
As Table˜7 shows, probes achieve very high F1 in the detection of synthetically created hallucinations across all tasks. However, probe performance is subject to a significant drop when trained and tested on different modalities, with probes generally performing better when trained on organic hallucinations and tested on synthetic ungrounded responses compared to the reverse, with the notable exception being abstractive summarization on the CDM dataset. Prior work has shown that synthetic data generation approaches specifically designed for factuality evaluation Ma et al. (2023) do not align with actual errors made by generation models in their lexical error distributions Goyal and Durrett (2021). Here, our results show this to extend to their hidden state signals as well, which are used by the probe to detect unfaithfulness in the case of forced decoding on human-created content, relative to organic instances of hallucination.
Task Train Test F1-R
CDM Organic Organic 0.45
Synthetic Organic 0.48
Organic Synthetic 0.38
Synthetic Synthetic 0.93
CF Organic Organic 0.88
Synthetic Organic 0.65
Organic Synthetic 0.83
Synthetic Synthetic 0.94
E2E Organic Organic 0.87
Synthetic Organic 0.69
Organic Synthetic 0.90
Synthetic Synthetic 0.98
Table 7: Response-level PoolingE probe performance, as measured by F1-R, in detecting hallucinations on organic or synthetic ungrounded responses. Performance is generally better when the training data resembles the test data.
Task.
To test whether the saliency of grounding behavior persists consistently in similar hidden state locations across different tasks, we train our probe on each specific task and evaluate its performance when applied to the others. As shown in Table˜8 and compared to same-task training, we find a lack of generalization in hallucination detection performance for our probes across different train and test tasks. This may mean that there is not a consistent signal of hallucination across tasks, or that our probes are not distinguishing this signal from its task-or-dataset-specific confounds, and are overfitting to the confounds.
6 Discussion
In this work, we examined how LLM internal representations reveal their hallucinatory behavior, via a case study of several in-context grounded generation tasks. Our analyses point to several nuances in the saliency of this behavior across model layers, hidden state types, model sizes, hallucination types, and contexts. Below, we highlight a few takeaways as well as limitations that naturally suggest directions for future work.
Efficiency and Access.
Our work is motivated by the need for efficient LLM hallucination evaluation. Though computationally efficient, one of the key limitations of probing is in the need for labeled in-domain data for probe training. Our results in Table˜8 suggest poor generalization to out-of-domain tasks, although perhaps this would be mitigated by larger and more diverse training sets. Furthermore, it perhaps goes without saying that probing requires access to the hidden states themselves. If LLMs continue to move behind closed-source APIs, such hidden state probes will not be possible for those without privileged access.
Test Train F1-R
CDM CF 0.44
E2E 0.44
CF+E2E 0.46
CDM 0.45
CF CDM 0.60
E2E 0.80
CDM+E2E 0.75
CF 0.88
E2E CDM 0.60
CF 0.75
CDM+CF 0.76
E2E 0.87
Table 8: Response-level PoolingE cross-task probe performance, as measured by F1-R. All experiments use the same amount of training data (a random subset of the full training set). The "+" sign indicates an equal mixture of two training tasks.
Ecological Validity.
The conclusions from an ecologically valid experiment should hold outside the context of that study De Vries et al. (2020). Synthetic hallucinations are generally ecologically invalid for organic hallucination detection. Extending beyond surface-level error distributions Goyal and Durrett (2021), these differences in internal representations further suggest that language model behavior during hallucination fundamentally misaligns with current expectations of ungrounded generations. Training on the same modality—synthetic or organic—is imperative to maximize hallucination detection performance, and our results on the synthetic modality, especially on CNN / DailyMail, suggest probing to be a feasible method for complex multi-sentence entailment classification that can leverage the capability of LLMs.
Annotator Disagreements.
Despite asking annotators to label what they deem to be the minimal spans of hallucination in a response, considerable variation still exists between annotators on judgments of their exact span boundaries. It remains an open questions how to improve annotator agreement in this context.
Probe Design.
Even if span boundaries were standardized, they may not align perfectly with hallucination signals in LLM internal representations. For instance, while a generation may first contradict a given knowledge source at a certain token, the ultimate signal that it will hallucinate may already be salient in the tokens preceding it. Alternatively, the signal might be computed only later, as the decoder encodes previously generated tokens and recognizes that it is currently pursuing an ungrounded train of thought. Thus, there may be probe architectures that could improve hallucination detection performance and mitigation. Nonlinear probes such as feed-forward networks are also worth considering: the hidden layer(s) of a feed-forward network could learn to detect different types of hallucinations.
Mitigation.
Finally, once we can detect ungrounded generation behavior, we can try to learn to avoid it. This has great practical importance as retrieval-oriented generation becomes popular across a variety of consumer-facing applications.
We have left this subject for future work. A simple method is to reject a response or prefix where hallucinations have been detected with high probability, and start sampling the response again Guerreiro et al. (2023). Another approach would be to use the predicted hallucination probabilities to reweight or filter hypotheses during beam search. Finally, a detector can also be used as a proxy reward method for fine-tuning the language model, using a reward-based method such as PPO Schulman et al. (2017). In addition to learning from response-level negative rewards incurred at the end of the utterance, PPO can benefit from early token-level negative rewards, which serve as a more informative signal for reinforcement learning (i.e., reward shaping) and thus can reduce sample complexity.
Acknowledgments
We thank Val Ramirez, Patrick Xia, Harsh Jhamtani, Hao Fang, Tongfei Chen, and Radhika Gaonkar for their helpful comments, thoughts, and discussions, as well as the data specialists who contributed and aided in the creation of this work.
Appendix
Here, we provide additional information about annotation guidelines, annotator disagreement examples, prompts used to generate organic responses to our tasks, and ChatGPT baseline comparisons.
Appendix A Annotation Guidelines
The purpose of this annotation task is two-fold: (1) identify if a generated response is a hallucination relative to a given knowledge source, and (2) if so, identify and mark the location(s) of the minimal spans of text that are ungrounded against the knowledge source. A response is grounded in a knowledge source if that source entails the content of the response. In other words, the response should (1) not contradict or (2) draw conclusions that are not explicitly given in the knowledge source.
A.1 CNN / DailyMail
Provided Information.
(1) Original article, (2) summary of article.
Task.
Identify the presence and location of hallucinations in the summary of the article, relative to the original article text. You are to only annotate—by copying the text in the summary and adding characters < and > in specific locations—around the minimal span(s) of the summary text relative to the original article; I.e., if changing an additional token wouldn’t change whether or not the response was grounded or not, don’t include it in the span you annotate. Note: You will be given summaries that contain a variety number of sentences. You are only to annotate the first three sentences in the summary, discarding the rest.
Example.
(CNN) – The world’s fastest man on a pair of skis just got faster. On Monday Italian Simone Origone broke his own speed skiing world record as he reached 252.4 kilometers per hour on the Chabrieres slopes in the French Alps – an achievement confirmed by organizers France Ski de Vitesse. With a 1,220 m slope that has a maximum gradient of 98% and an average of 52.5%, Chabrieres is not for the faint-hearted. Traveling at over 250 km/h is a privilege usually reserved for Formula One drivers, yet speed skier Origone was equipped with just an aerodynamic helmet to increase streamlining and a ski suit made from air-tight latex to reduce wind resistance. Origone’s one nod to safety was wearing a back protector in case of a crash as he threw himself down the one kilometer track. Origone has been the fastest speed skier on the globe since April 2006, having set a then-new world record of 251.4 km/h at Les Arcs. Bastien Montes of France was Origone’s closest challenger on Monday, but even the Frenhcman’s new personal best of 248.105 km/h was someway short of the Italian’s 2006 world record, let alone his latest one. "Simone Origone is the greatest champion of all time, he is the only person to hold the record for two ski speeds in France – Les Arcs and Vars Chabrieres. This is a historic day," technical director of Speed Masters Philippe Billy told Vars.com. The 34-year-old Origone – a ski instructor, mountain guide and rescuer by day – only took up the discipline in 2003, having given up downhill skiing in 1999. "Now that I have twice won the world record, I can say that I have made history," Origone told Vars.com. "It is important for me and for speed skiing."
Hypothetical Instance 1.
The world’s fastest man on skis just got faster.
  • Everything is grounded, no need to mark anything here.
Annotation. The world’s fastest man on skis just got faster.
Hypothetical Instance 2.
The world’s fastest man on skis just got faster. He has now won the world record twice. He has excelled at Chabrieres. He has been the fastest speed skier on the globe since April 2006. His speed record reached more than 260 kilometers per hour.
  • This summary has more than 3 sentences, so the first thing to do is to only look at the first three and discard any sentences beyond that in your annotation.
  • Even though the last sentence is ungrounded relative to the source text, it is not within the first three sentences, so we can ignore it here.
  • Annotate as such, copying the first three sentences into the annotation box. No need to highlight anything, as everything in the first three sentences are grounded:
Annotation. The world’s fastest man on skis just got faster. He has now won the world record twice. He has excelled at Chabrieres.
Hypothetical Instance 3.
Simone Origone, now in his twenties, just got faster. He has now won the world record more than twice.
  • “In his twenties” contradicts with “The 34-year-old Origone” in the original article.
  • “more than twice” contradicts with how he has now only won it exactly twice. As such, annotate:
Annotation. Simone Origone, now in his <twenties>, just got faster. He has now won the world record <more than> twice.
Special Points of Note.
Below, we list a few points of special note to keep in mind during your annotations.
Concatenation.
When finding yourself annotating something like “The city’s pollution levels have reached or <exceeded ><“very high”> levels for the last…,” do concatenate the two labeled spans; if there are no characters between them (for example, <exceeded ><“very high”>), concatenating the annotation bounds into (<exceeded “very high”>) is ideal.
“Was”/“Is”.
If the article says something like “Woods insisted he is ready to win the famous Claret…” and the summary said something like “said he was ready to win the famous Claret…”, “was” here would not be considered to be a hallucination, as the summary is recounting what was said in the article, so the past tense here doesn’t contradict what was originally said in the article.
When The Article Has “Contradicting” Facts.
Articles may sometimes “correct” themselves; i.e. the article may have something where it originally states a claim but then corrects it later in the article. For example, “The NWS – last May – initially forecast 13 to 20 “named storms,” including seven to 11 hurricanes. Then in August, it dialed down that initial forecast to predict 13 to 19 named storms, including six to nine hurricanes.”
In this case, if the summary only states “The National Weather Service forecast called for 13 to 20 named storms” (and doesn’t mention it was corrected later to 13 to 19), then this is a hallucination as it is corrected later in the article.
If the summary states “The National Weather Service forecast initially called for 13 to 20 named storms” (and doesn’t mention it was corrected later to 13 to 19) then this is not a hallucination, as it quantifies the original prediction with “initially.”
If the summary states “The National Weather Service forecast called for 13 to 19 named storms,” then this is not a hallucination, as it is the final corrected fact.
Identifying the Summary.
Sometimes the summarization system may produce things that are not a summary of the article. For example, “Woman dies while paragliding in Tenerife. Paragliding in Tenerife. Paragliding in Tenerife. Paragliding in Tenerife. Paragliding Tenerife Canary Islands.”
Here, only the first sentence is the summary, and you can discard the rest. This is because the summarization system tries to mimic how actual website summaries work, so they might also add things like topic tags, keywords, and things like that.
Another example is “InvisiBra, £48, Lavalia, lavalia.co.uk; Strapless Stick On Adhesive Bra, £12, Marks and Spencer; Natural Stick On Bra, £20, Debenhams.” These are all just keyword tags; no summary of the article is present here. Try to use your best judgment here, as it might not always be clear.
In the case where such things are generated that do not look to be part of a summary of the article in terms of format (not content), discarding them and everything after them is the way to go. For the former example, you would annotate only “Woman dies while paragliding in Tenerife.” for hallucinations.
A.2 Conv-FEVER
Provided Information.
(1) Conversational context, (2) Knowledge for next response, (3) Next response.
Task.
Identify the presence and location of hallucinations in “Next response”, relative to “Knowledge for next response”. You are to only annotate—by copying the text in “Next Reply” and adding characters < and > in specific locations—around the minimal span(s) of an ungrounded response relative to the knowledge; I.e., if changing an additional token wouldn’t change whether or not the response was grounded or not, don’t include it in the span you annotate.
Example.
Conversational Context: A: lets talk bowling a game where you throw a bowling ball towards a target. B: ok I love bowling I would like to know more A: for one it is one of the major forms of throwing sports and I think a lot of people enjoy it B: yes me too I use to bowl all the time with my family what about the strike? A: well a strike is when you knock down all pins on the first round B: ok how can I get a turkey I think that may be hard to get though Knowledge for Next Reply: Three consecutive strikes are known as a "turkey".
Hypothetical Instance 1.
Next Reply: A: yea four consecutive strikes is known as a turkey
  • Here, a contradiction is present in the underlined portions: four consecutive strikes in the response goes against what is provided in the knowledge, which says three.
  • To make this reply grounded, one would have to change “four” to “three”.
  • As such, in your annotation, you will highlight the minimum needed span of text to make this reply grounded, with chevrons <>. Here, the minimum needed span of text to change is “four”, so you would annotate as such:
Annotation. yea <four> consecutive strikes is known as a turkey
Hypothetical Instance 2.
Next Reply: A: yea it is pretty hard, four consecutive strikes is known as a turkey
  • Here, the contradiction is still present in the previously underlined portion: four is a contradiction with three.
  • However, in addition to this contradiction, you see an expression of an opinion in this response: that getting a turkey is pretty hard (underlined).
  • You are to ignore the expression of personal opinions during your annotation of minimal spans for this task, as we won’t consider conversational chitchat—if complemented with grounded information elsewhere in the reply—to be a hallucination for this task. As such, you would annotate this as:
Annotation. yea it is pretty hard, <four> consecutive strikes is known as a turkey
Hypothetical Instance 3.
Next Reply: A: yea it is pretty hard.
  • Here, the response is purely an expression of an opinion and does not use any information in the provided knowledge.
  • As such, you would mark this entire response to be a hallucination, annotating as:
Annotation. <yea it is pretty hard.>
Hypothetical Instance 4.
Next Reply: A: Three strikes consecutively is a turkey.
  • Here, the response is grounded. As such, you won’t mark anything in the response, annotating as such:
Annotation. Three strikes consecutively is a turkey.
Hypothetical Instance 5.
Next Reply: A: four consecutive strikes is known as a turkey, also known as a strike.
  • Two hallucinations are present in this response:
  • The direct contradiction of four with three in the knowledge source, and
  • The final part of the reply, saying that it’s also known as a strike, which is information not provided anywhere in the knowledge source. As such, you would mark this as:
Annotation. <four> consecutive strikes is known as a turkey, <also known as a strike>
A.3 E2E
Provided Information.
(1) Restaurant attributes, (2) Generated restaurant description.
Task.
Identify the presence and location of hallucinations in the generated restaurant description—by copying the text in the summary and adding characters < and > in specific locations—around the minimal span(s) of the generated restaurant description relative to the provided attributes. i.e., if changing an additional token wouldn’t change whether the description was grounded or not, don’t include it in the span you annotate. Note that numerical ratings should be rendered as numerical ratings and text ratings should be rendered as non-numerical expressions. For example, “customer rating” can be “low”, “medium” and “high”, but also “1”, “3”, and “5”. If the customer rating is “1”, for instance, and the generated description says “…has a low customer rating”, that would be considered a hallucination. The “priceRange” rating also has this attribute, in addition to “customer rating.” Finally, E2E specifies the "familyFriendly" attribute as akin to Child-friendly in gold annotations: if a restaurant is family-friendly, then it is also child-friendly.
Hypothetical Instance 1.
Restaurant Attributes: name[The Wrestlers], priceRange[less than £20], customer rating[low], familyFriendly[no]. Restaurant Description: The Wrestlers is an adults only restaurant with a customer rating of 1 out of 5 and a price range of less than £20.
  • While the familyFriendly[no] attribute says that the restaurant is not family friendly, this does not mean that the restaurant is adults only.
  • - Having a low customer rating does not imply a 1 out of 5 customer rating. As such, annotate as:
Annotation. The Wrestlers is an <adults only> restaurant with a customer rating of <1 out of 5> and a price range of less than £20.
Hypothetical Instance 2.
Restaurant Attributes: name[The Dumpling Tree], eatType[restaurant], food[Italian], priceRange[high]. Restaurant Description: The Dumpling Tree is an Italian restaurant with high prices.
  • Everything is grounded, no need to mark anything.
Annotation. The Dumpling Tree is an Italian restaurant with high prices.
Hypothetical Instance 3.
Restaurant Attributes: name[The Wrestlers], priceRange[less than £20], customer rating[low], familyFriendly[no]. Restaurant Description: Prices will never exceed £20 at The Wrestlers and they also don’t really rate well. They are not even family-friendly.
  • Having a priceRange attribute of less than 20 does not mean that prices will never exceed twenty.
  • As “will never exceed” is the minimal span you would need to change to make the description grounded relative to the provided attributes, annotate:
Annotation. Prices <will never exceed> £20 at The Wrestlers and they also don’t really rate well. They are not even family-friendly.
Hypothetical Instance 4.
Restaurant Attributes: name[Midsummer House], food[French], customer rating[5 out of 5], near[Café Rouge]. Restaurant Description: A place for fine dining French food, Midsummer House is nearby Cafe Rouge.
  • Nowhere in the attributes does it say that it’s “fine dining.” As such, annotate:
Annotation. A place for <fine dining> French food, Midsummer House is nearby Cafe Rouge.
Hypothetical Instance 5.
Restaurant Attributes: name[Browns Cambridge], priceRange[cheap], priceRange[moderate], customer rating[1 out of 5]. Restaurant Description: The Brown’s chain of food restaurants is an 1 out of 5 customer rated restaurant, and have moderate and cheap pricing. They are childfree.
  • The attribute does not mention that Browns Cambridge is a chain of food restaurants.
  • Furthermore, it does not say that the restaurant is childfree. As such, annotate:
Annotation. The Brown’s <chain of food restaurants> is an 1 out of 5 customer rated restaurant, and have moderate and cheap pricing. <They are childfree.>
Hypothetical Instance 6.
Restaurant Attributes: name[Browns Cambridge], priceRange[more than £30]. Restaurant Description: The Browns Cambridge is a restaurant with high prices.
  • More than £30 does not necessarily mean “high prices.” As such, annotate:
Annotation. The Browns Cambridge is a restaurant with <high prices>.
Taxi driver dropped off a soldier on an unsafe road, and he was killed by a coach. The taxi driver was not at fault.
Taxi driver dropped off a soldier on an unsafe road, and he was killed by a coach. The taxi driver was not at fault.
Accounting is the language of business, and it measures the results of an organization’s economic activities.
Accounting is the language of business, and it measures the results of an organization’s economic activities.
Alimentum is a high rated family-unfriendly restaurant with less than £20 per person.
Alimentum is a high rated family-unfriendly restaurant with less than £20 per person.
Table 9: Annotator span-level annotation differences for hallucination, annotated in bold, in organic model responses for CNN / DailyMail (top), Conv-FEVER (middle), and E2E (bottom).
Appendix B Annotator Disagreements
Examples of annotator disagreements in the exact spans of hallucination are shown in Table˜9. Qualitatively, we find that annotators agreed generally on the specific items of hallucination; i.e. the subject that was hallucinated (i.e. the taxi driver), the description of what was mentioned (i.e. the description of accounting), and the attribute that was hallucinated (i.e. the specific rating of the resturant). However, where exactly those hallucinated spans began and ended was a source of reasonable variation among annotators.
Appendix C Prompts
Organic Model Responses.
CNN / DailyMail generations take the form of a 0-shot prompt with [ARTICLE] TL;DR: [SUMMARY], whereafter the first 3 sentences generated in the summary are taken to be the summary of the article. For Conv-FEVER, we use a 2-shot in-context learning prompt in the form
## CONVERSATION HISTORY [CONTEXT]## KNOWLEDGE FOR NEXT RESPONSE[KNOWLEDGE]## RESPONSE THAT USES THE INFORMATION IN KNOWLEDGE[RESPONSE]
whereafter everything generated in the response preceding a newline character is taken as the final generation. For E2E, we use a 10-shot in-context learning prompt in the form of
Restaurant Attributes: [RESTAURANT ATTRIBUTES]Restaurant Description: [RESTAURANT DESCRIPTION]
with the same newline constraints as Conv-FEVER.
C.1 ChatGPT Prompts
ChatGPT Classic.
For the most basic zero-shot prompt classifier with no reasoning, we use the following prompt for CNN / DailyMail, following Luo et al. (2023):
Decide if the following summary is consistent with the corresponding article. Note that consistency means all information in the summary is supported by the article.Article: [Article]Summary: [Summary]Answer (yes or no):
For Conv-FEVER, our prompt is structured as follows:
Decide if the following sentence (SENTENCE) is consistent with the corresponding knowledge statement (KNOWLEDGE). Note that consistency means all information in the sentence (SENTENCE) is supported by the knowledge statement (KNOWLEDGE).Knowledge Statement (KNOWLEDGE): [Knowledge]Sentence (SENTENCE): [Response]Answer (yes or no):
Finally, for E2E, our prompt is:
Decide if the following restaurant description (DESCRIPTION) is consistent with the provided restaurant attributes (ATTRIBUTES). Note that consistency means all information in the description is supported by the provided attributes.Restaurant Attributes (ATTRIBUTES): [Attributes]Restaurant Description (DESCRIPTION): [Response]Answer (yes or no):
ChatGPT CoT.
For zero-shot chain-of-thought reasoning prompts, we simply replace Answer (yes or no): in all prompts with Explain your reasoning step by step then answer (yes or no) the question:, following Luo et al. (2023).